OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = 3x2 + 24x - 1?

OpenStudy (dumbcow):

for y = ax^2 +bx+c line of symmetry = -b/(2a)

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

i have answer choises

OpenStudy (anonymous):

x = -8 x = -6 x = -4

OpenStudy (welshfella):

compare your equation with the general form and you'll be able to see the values of a,b and c y = ax^2 + bx + c y = 3x^2 + 24x - 1

OpenStudy (welshfella):

can you see what a and b are?

OpenStudy (anonymous):

no... :(

OpenStudy (welshfella):

what number lies directly below a ?

OpenStudy (welshfella):